OpenStudy (anonymous):

sin^2(theta)+cos(theta)=2 Use the Pythagorean identity sin^2(theta)+cos^2(theta)=1 to replace sin^2(theta) in the given equation

OpenStudy (mertsj):

Let's use x instead of theta.

OpenStudy (mertsj):

\[\sin ^2x+\cos x=2\] \[1-\cos ^2x+\cos x=2\]

OpenStudy (mertsj):

\[\cos ^2x-\cos x+1=0\]